Circular No. 7816

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V838 MONOCEROTIS S. Kimeswenger, C. Lederle, and S. Schmeja, University of Innsbruck,

report that this object has increased markedly in brightness, as seen by their photometry (cf. IAUC 7796): Feb. 1.8576 UT, B = 12.794, V = 10.708, R = 9.652, I_{c} = 8.634; 2.7986, 9.607, 8.193, 7.439, 6.657; 2.9132, 9.427, 8.024, 7.276, 6.484. Spectroscopic follow-up is encouraged.

Visual magnitude estimates: Jan. 19.137 UT, 10.3 (R. Y. Shida, São Paulo, Brazil); 26.798, 10.8 (A. Pearce, Nedlands, W. Australia); Feb. 2.803, 8.3 (R. J. Bouma, Groningen, The Netherlands); 2.847, 8.9 (M. Reszelski, Szamotuly, Poland); 2.843, 8.6 (O. Pejcha, Brno, Czech Rep.).

SUPERNOVA 2002ap IN M74 S. Smartt and E. Ramirez-Ruiz, Institute of Astronomy, Cambridge;

and P. Vreeswijk, University of Amsterdam, report that 120-s UBV I images of M74 were taken with the Isaac Newton Telescope (+ Wide Field Camera) in July 2001, as part of the INT Wide Field Survey. They detect an object with B = 21.6, V = 21.2, I = 20.5 (uncertainties ± 0.2 mag) at α = 1^{h}36^{m}24^{s}.00, δ = +15^{o}45′13′′.6 (equinox 2000.0), which is within 1′′.3 of a radio source found on Feb. 1.03 UT (cf. http://gcn.gsfc.nasa.gov/gcn/ gcn3/1237.gcn3; position end figures 23^{s}.920, 12′′.87). This fairly faint object shows some inconclusive evidence of being extended; at the spatial resolution of the images (∼ 50 pc), the object could be an unresolved small cluster or H II region, or a predominantly bright star on a variable background. Assuming that it is a single star, with a distance modulus of 29.5 to M74, they estimate absolute magnitudes M_{B} = −8.2, M_{V} = −8.5, M_{I} = −9.1. The colors B −V = +0.3, V −I = +0.6, and these absolute magnitudes are consistent with a very luminous early- to mid-F-type supergiant. The bolometric luminosity is ≈ 10^{5.3} L_{⊙}, which would place the star in the luminous-blue-variable region of the HR diagram (Humphreys and Davidson 1994, PASP 106, 1025) and suggests an initial mass of around 40 M_{⊙}. Smartt et al. add that any significant extinction would make the star intrinsically bluer and more luminous. Further astrometry and image-shape analysis are required to confirm that the progenitor object is stellar, and monitoring of SN 2002ap at all wavelengths is encouraged.

Visual magnitude estimates: Jan. 31.784 UT, 13.2 (L. Brat, Pec pod Snezkou, Czech Republic); Feb. 2.776, 12.7 (R. J. Bouma, Groningen, The Netherlands); 2.783, 12.6 (K. Hornoch, Lelekovice, Czech Republic).

2002 February 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green